### v1

# SALAD: Self-Adaptive Link Adaptation

Abstract
Adapting the modulation and coding scheme (MCS) to the wireless link quality is critical for maximizing spectral efficiency while ensuring reliability. We propose SALAD (self-adaptive link adaptation), an algorithm that exclusively leverages ACK/NACK feedback to reliably track the evolution of the signal-to-interference-plus-noise ratio (SINR), achieving high spectral efficiency while keeping the long-term block error rate (BLER) near a desired target. SALAD infers the SINR by minimizing the cross-entropy loss between received ACK/NACKs and predicted BLER values, with a learning rate that self-adapts online through knowledge distillation. Based on this inference, SALAD selects the MCS via hypothesis testing: if the SINR is likely underestimated, a higher MCS is selected to accelerate link adaptation under improving channel conditions. To prevent BLER drift from its long-term target, SALAD incorporates a feedback control loop that adjusts the instantaneous BLER target. Over-the-air experiments on a 5G testbed demonstrate that SALAD consistently outperforms the industry-standard outer-loop link adaptation (OLLA). With a single set of parameters, SALAD achieves up to 15% higher throughput and spectral efficiency than multiple OLLA variants across different traffic regimes, while meeting the BLER target.
I Introduction
Adapting the modulation and coding scheme (MCS) to varying wireless channel conditions, known as link adaptation (LA), is essential for achieving high spectral efficiency (SE). Higher SE translates into increased user throughput, lower latency, and reduced transmission energy thanks to shorter transmission durations and fewer retransmissions. An effective LA strategy must also guarantee a sufficiently low block error rate (BLER), as excessive re-transmissions waste resources, increase latency, and degrade performance. In fact, transport protocols such as TCP misinterpret packet losses caused by link-level errors as signs of network congestion, resulting in an unwarranted reduction in transmission rate [1]. For these reasons, a target BLER value is typically defined to strike a balance between throughput and reliability. Empirical evidence and theoretical analyses [2] support a fixed target BLER value of approximately $10\%$, although adapting the target dynamically based on the current signal-to-interference-plus-noise ratio (SINR) can improve throughput [3].
LA techniques can be broadly categorized into sampling-based and model-based approaches. Sampling-based methods, such as auto rate fallback (ARF) [4] and sample rate [5], adapt the MCS by trial-and-error, solely relying on statistics on transmission successes (ACK) and failures (NACK) for each available configuration, received via HARQ (hybrid automatic repeat request) feedback. Due to their slow convergence, these methods are mainly suitable for scenarios with few MCS configurations and relatively static channels, as in Wi-Fi.
In contrast, model-based methods leverage statistical models of the wireless channel and side-information, such as SINR estimates, to make more informed decisions. A prominent example is outer-loop link adaptation (OLLA) [6], the current industry-standard for cellular networks. OLLA selects the MCS that meets a target BLER according to pre-computed BLER tables and the latest SINR estimation. The latter is adjusted by an offset which is updated according to a fixed stepsize rule based on the received ACK/NACKs.
Despite its simplicity, OLLA is grounded in stochastic approximation theory (see Section II) and can provably drive the long-term BLER toward the desired target. Although OLLA’s fixed stepsize works reasonably well across different mobility scenarios, it excels in none. A small stepsize results in slow adaptation to rapidly changing channel conditions, which is particularly detrimental for short-lived connections [7]. Conversely, a large stepsize can cause instability in static regimes, leading to SE loss and/or short-term BLER degradation. Our contribution aims at combining the best of both worlds: the stepsize for SINR updates increases as the channel changes more rapidly.
SINR estimation is a notoriously difficult challenge, especially for the base station (BS) in the downlink (DL), which is our main focus. Considering 5G NR with time-division multiplexing (TDD), channel reciprocity allows the BS to estimate the useful signal strength via sounding reference signals (SRS), although interference remains unobservable. An alternative common approach is to let the user equipment (UE) estimate the SINR based on channel state information reference signals (CSI-RS) and report it to the BS using the channel quality indicator (CQI). Conversely, in frequency-division multiplexing (FDD), the only option is to rely on CQI reports [8]. Yet, CQI reports present several well-known limitations. First, their granularity is coarse, as the CQI is an integer ranging from 0 to $15$. Second, the reports can be infrequent, with periodicity configurable from a few to several hundreds of slots. Moreover, even when frequent, CQI reports are still delayed in TDD systems, as they can only be transmitted during uplink slots. Finally, reports are wide-band, while transmission often occurs over a narrow band. Motivated by this, we reduce the reliance on CQI reports by introducing a principled method to infer the SINR using only ACK/NACK feedback.
I-A Problem formulation
To formalize the link adaptation problem, we consider a single user and denote its effective [^1] The effective SINR, or simply called SINR throughout the text, of a user scheduled on different time-frequency resources is defined as the SNR of an equivalent AWGN channel that yields the same BLER as the original channel. signal-to-interference-plus-noise ratio (SINR) by $\gamma_{t}$. We consider that the transport block size (TBS) $b_{t}$ is provided by the MAC scheduler. [^2] In practice, the MCS selection may occur prior to TBS allocation. In this case, the TBS can be estimated from past allocations. At each slot $t$, the MCS index $u_{t}$ is ideally selected to maximize the expected SE, defined as the SE [^3] The SE is defined as the modulation order, i.e., the number of bits per symbol, multiplied by the code rate. of the selected MCS $u$ multiplied by the probability of successful transmission, while not exceeding the predefined BLER target $\tau\in(0,1)$:
$\max_{u\in U}\,\mathrm{SE}(u)\times\left(1-\text{\small{BLER}}(u,\gamma_{t},b_{t})\right)$ (1)
$\ \mathrm{s.t.}\ \text{\small{BLER}}(u,\gamma_{t},b_{t})\leq\tau,$
where $\mathrm{SE}(u)$ is the SE corresponding to MCS $u$ and $U$ is the discrete set of available MCS indices.
If the SINR $\gamma_{t}$ is known perfectly, then (1) can be solved by brute force search. Indeed, the MCS’s SE is known, while the BLER can be reliably pre-computed via simulations on an additive white Gaussian noise (AWGN) channel, for various combinations of SE, SINR, and TBS. In practice, however, the true SINR is unknown and can only be estimated, with accuracy varying across scenarios, as discussed in Section I. As a result, link adaptation requires a non-trivial joint inference of the unknown SINR and control of the selected MCS at each slot.
I-B Our contribution
We propose self-adaptive link adaptation (SALAD), a novel algorithm that estimates the SINR by minimizing the cross-entropy loss between the observed ACK/NACK feedback and the BLER via gradient descent, as described in Section III-A. Its learning rate can be adjusted in an online fashion by a mechanism inspired by knowledge distillation, which compares the gradient-based estimate to a more powerful batch estimator. When the SINR is deemed to be underestimated—based on a hypothesis testing—SALAD increases the instantaneous BLER target to probe a higher MCS value and corrects the SINR estimation accordingly (Section III-B). To meet a desired long-term BLER, SALAD adjusts the instantaneous BLER target via a feedback loop (Section III-C).
We demonstrate that SALAD outperforms OLLA in system-level simulations and over-the-air experiments conducted within a controlled setup comprising a rotational absorber table between a commercial off-the-shelf (COTS) open-RAN radio unit (O-RU) and a COTS user equipment (UE) in an anechoic chamber. In Section IV, we show that, with a single set of pre-defined parameters, SALAD achieves up to 1515% higher throughput and SE compared to multiple variants of OLLA with different stepsizes, all while reliably meeting the BLER target. Moreover, SALAD is more robust to delayed CQI reports than OLLA. Finally, parameter tuning allows SALAD to adapt to extreme throughput scenarios, yielding up to an additional 33% throughput gain.
The code used to produce the results presented in Section IV-C is publicly available at [9].
TABLE I: List of symbols.
Symbol Meaning
$t$ Upcoming slot index
$\gamma$ True signal-to-interference-plus-noise ratio (SINR)
$\gamma^{\mathrm{rep}}$ SINR corresponding to the last CQI report
$\text{\small{BLER}}^{\mathrm{tab}}$ Block error rate (BLER) from pre-computed table
$\tau$ Long-term BLER target
$1{\thinspace}\mathrm{I}^{\textsc{NACK}}$ NACK indicator (1 if NACK, 0 if ACK)
$u$ Modulation and coding scheme (MCS) index
$\mathrm{SE}(u)$ Spectral efficiency of MCS $u$
$b$ Transport block size (TBS)
$\delta$ Number of in-flight packets
$\sigma(.)$ Logistic sigmoid function
$c,s$ Sigmoid center/scale for BLER approximation
$\widetilde{\gamma}^{\mathrm{OLLA}}$ OLLA’s estimated SINR
$\Delta$ OLLA’s SINR offset
$\Delta^{\textsc{NACK}}$ OLLA’s offset stepsize adjustment upon a NACK
$\Delta^{\textsc{ACK}}$ OLLA’s offset stepsize adjustment upon an ACK
$\widetilde{\gamma}$ SALAD’s estimated SINR
$\varepsilon$ SALAD’s student model learning rate
$\theta$ SALAD’s teacher model parameters ($\theta\in\mathbb{R}^{K}$)
$\beta$ SALAD’s teacher model regularization coefficient
$E$ SALAD’s integral BLER error
$k_{E}$ SALAD’s integral BLER error coefficient
$\tau_{t}$ SALAD’s instantaneous BLER target at slot $t$
$\tau^{\mathrm{probe}}$ SALAD’s high BLER target for MCS probing
$\mathcal{H}_{0}$ Null hypothesis that SINR estimates are correct
$\mathcal{S}$ SALAD’s bias score
$\rho$ SALAD’s bias score ratio threshold for MCS probing
$T$ SALAD’s bias score window length (n. samples)
$p^{\mathrm{probe}}$ SALAD’s probing probability if score ratio $\rho$
I-C Related work
In the context of cellular networks, OLLA was first proposed for link adaptation in the early 2000s [10, 6], although its mechanism and theoretical justification via Markov chain arguments dates back to the 1990s, initially applied to power control [11]. Several heuristic extensions of OLLA have since been proposed. These include stepsize adaptation [12, 13], offset dynamics optimization via reinforcement learning (RL) [14], initialization strategies of the offset based on prior connection statistics [15], parameter tuning through deep learning [16], enhanced feedback mechanisms for ultra-reliable connections [17], and customization for emerging applications such as extended reality [18]. Alternative link adaptation strategies with firm theoretical foundations formulate the problem as a multi-armed bandit (MAB), where each MCS corresponds to an “arm” [19, 20]. However, these methods often exhibit slow convergence, similar to other sampling-based methods originally designed for Wi-Fi, such as ARF [4] and sample rate [5]. To address this limitation, [21] formulates a contextual MAB framework that jointly learns the BLER for each of the MCS configuration given the current network context. In [22], a Bayesian approach is used to learn the BLER for each MCS, with Thompson sampling guiding the MCS selection process. Subsequently, [23] proposes latent Thompson sampling to infer the underlying SINR. However, such approaches are not designed to accommodate rapid SINR variations and do not explicitly target a predefined BLER. A supervised learning framework is introduced in [24] to directly predict the BLER. A practical approach that combines deep learning for BLER prediction and a calibration method to compensate for prediction errors is studied in [25]. The work in [26] employs offline RL approach via decision transformers under the assumption of persistent CQI reporting.
II The industry standard: OLLA
Since the adoption of 3G, the industry standard for link adaptation in cellular systems is the outer-loop link adaptation (OLLA) algorithm [6]. To drive the long-term BLER close to the target value $\tau$, OLLA maintains an additive offset variable $\Delta$ [ $\text{\,}\mathrm{dB}\text{/}$] that corrects the latest SINR estimate $\gamma^{\mathrm{rep}}$. The latter can be obtained by CQI reports as explained in [8]. When a NACK is received via the HARQ feedback (denoted by the indicator function $1{\thinspace}\mathrm{I}^{\textsc{NACK}}=1$), the offset is decreased by a predefined value $\Delta^{\textsc{NACK}}$. Conversely, an ACK ($1{\thinspace}\mathrm{I}^{\textsc{NACK}}=0$) leads to a offset increase of $\Delta^{\textsc{ACK}}$, computed as:
$\Delta^{\textsc{ACK}}=\frac{\tau}{1-\tau}\Delta^{\textsc{NACK}}.$ (2)
We now fix slot $t$. Let $\delta_{t}$ be the number of in-flight packets at slot $t$, for which the HARQ feedback has not been received yet. Then, the latest SINR estimate available at slot $t$ is $\widetilde{\gamma}_{t-\delta_{t}}$, computed as in line 12 of Algorithm 1. Then, OLLA assigns to the user the largest MCS whose BLER does not exceed the target $\tau$, according to the latest SINR estimate:
$u^{\mathrm{ILLA}}($ $\widetilde{\gamma}^{\mathrm{OLLA}}_{t-\delta_{t}},\tau,b_{t}):=$
$\operatorname{argmax}\big\{u\!:\text{\small{BLER}}^{\mathrm{tab}}\!\left(u,\widetilde{\gamma}^{\mathrm{OLLA}}_{t-\delta_{t}},b_{t}\right)\leq\tau\big\},$ (3)
where $\text{\small{BLER}}^{\mathrm{tab}}$ contains pre-computed BLER values, simulated on the AWGN channel. This step is commonly called inner-loop link adaptation (ILLA) [^4] An alternative, drop-in replacement for ILLA selects the MCS that maximizes the expected spectral efficiency in (3)..
Algorithm 1 Outer-Loop Link Adaptation (OLLA), e.g., [6]
1:Parameters BLER target $\tau$, NACK stepsize $\Delta^{\textsc{NACK}}$.
2:Set ACK offset update $\Delta^{\textsc{ACK}}=\tfrac{\tau}{1-\tau}\Delta^{\textsc{NACK}}$
3:Initialize $\Delta:=0$, $\gamma^{\mathrm{rep}}:=0$
4:for slot $t=0,1,\dots$ do
5:  Receive set $\mathcal{N}$ of ACK/NACK and CQI report (if any).
6:  if CQI is reported then
7:   Convert it to SINR $\gamma^{\mathrm{rep}}$ [ $\text{\,}\mathrm{dB}\text{/}$]
8:  end if
9:  for $1{\thinspace}\mathrm{I}^{\textsc{NACK}}\in\mathcal{N}$ do
10:   Update the SINR offset as:
$\Delta\leftarrow\left\{\begin{array}[]{ll}\!\!\!\Delta+\Delta^{\textsc{ACK}}&\text{if }1{\thinspace}\mathrm{I}^{\textsc{NACK}}=0\ \ \text{(ACK)}\\ \!\!\!\Delta-\Delta^{\textsc{NACK}}&\text{if }1{\thinspace}\mathrm{I}^{\textsc{NACK}}=1\ \ \text{(NACK)}\end{array}\right$. (4)
11:  end for
12:  Set the latest SINR estimate to $\widetilde{\gamma}^{\mathrm{OLLA}}_{t-\delta_{t}}\!:=\!\gamma^{\mathrm{rep}}\!+\!\Delta$ [ $\text{\,}\mathrm{dB}\text{/}$]
13:  if user is scheduled then
14:   Select the MCS as $u^{\mathrm{ILLA}}(\widetilde{\gamma}^{\mathrm{OLLA}}_{t-\delta_{t}},\tau,b_{t})$
15:  end if
16:end for
II-A OLLA through the lens of stochastic approximation
To build a solid foundation for our algorithm, it is instructive to reinterpret OLLA within the context of stochastic approximation (SA). Consider an unknown non-decreasing function $f$ and a constant target $\tau$, and assume that the equation $f(y)=\tau$ has a root at $y=y^{*}$ that we want to find. The function $f$ can only be observed via noisy samples $\widetilde{f}$, providing an unbiased estimate of the true value, i.e., $\mathbb{E}[\widetilde{f}(y)]=f(y)$ for all $y$.
SA starts with an initial guess $y_{0}$ of the root $y^{*}$; at time $t>0$, upon observing $\widetilde{f}(y_{t-1})$, it updates its guess as [27]:
$y_{t}=y_{t-1}+\alpha_{t}\left(\tau-\widetilde{f}(y_{t-1})\right),$ (5)
where $\alpha_{t}>0$ is SA’s stepsize. If $\widetilde{f}(y_{t-1})$ undershoots the target $\tau$, i.e., $\widetilde{f}(y_{t-1})<\tau$, then $y_{t-1}$ is likely an underestimation of the root $y^{*}$, since $f$ is non-decreasing. Thus, $y_{t}>y_{t-1}$, and their difference $y_{t}-y_{t-1}$ is proportional to the undershoot. The reverse argument holds in case of an overshoot.
OLLA is indeed an SA algorithm under two assumptions:
  • i)
    the user’s SINR report is fixed, e.g., $\gamma^{\mathrm{rep}}:=0$
  • ii)
    the ACK/NACK feedback $1{\thinspace}\mathrm{I}^{\textsc{NACK}}_{t}$ is received immediately upon MCS selection $u_{t}$, i.e., $\delta_{t}=0$ for all $t$.
In this case, the SINR estimate $\widetilde{\gamma}^{\mathrm{OLLA}}_{t}$ plays the role of $y_{t}$, and the unknown function $f$ at time $t$ is
$f_{t}(\widetilde{\gamma}^{\mathrm{OLLA}}_{t}):=\text{\small{BLER}}\left(u^{\mathrm{ILLA}}(\widetilde{\gamma}^{\mathrm{OLLA}}_{t},\tau,b_{t}),\gamma_{t},b_{t}\right),$ (6)
which is time-varying due to the true SINR $\gamma_{t}$ and the TBS $b_{t}$. The function $f_{t}$ is observed via the feedback $1{\thinspace}\mathrm{I}^{\textsc{NACK}}_{t}$, which plays the role of the noisy observation $\widetilde{f}_{t}(\widetilde{\gamma}^{\mathrm{OLLA}}_{t})$. The latter is indeed unbiased since $\mathbb{E}[1{\thinspace}\mathrm{I}^{\textsc{NACK}}_{t}]=f_{t}(\widetilde{\gamma}^{\mathrm{OLLA}}_{t})$.
The parallel between SA and OLLA becomes clear once we rewrite the offset update (4) as:
$\widetilde{\gamma}^{\mathrm{OLLA}}_{t}=\widetilde{\gamma}^{\mathrm{OLLA}}_{t-1}+\frac{\Delta^{\textsc{NACK}}}{1-\tau}\left(\tau-1{\thinspace}\mathrm{I}^{\textsc{NACK}}_{t-1}\right)\quad\forall\,t\geq 0.$ (7)
In fact, (7) and (5) show the same structure, with the term $\frac{\Delta^{\textsc{NACK}}}{1-\tau}$ in (7) acting as the SA’s constant stepsize $\alpha$.
This parallel confirms that OLLA is effectively designed to achieve a predefined BLER target $\tau$. We proceed by discussing OLLA’s stepsize choice under the perspective of SA.
Figure 1: OLLA’s behavior when the SINR switches between two values and HARQ feedback delay $\delta=5$. For small values of $\Delta^{\textsc{NACK}}$, the MCS adapts slowly to the channel conditions, and the BLER (over a sliding window of 50 slots) exhibits high variance. For large $\Delta^{\textsc{NACK}}$, OLLA tracks the SINR variations more rapidly but introduces excessive MCS fluctuations in the stationary phase, causing SE degradation. The oracle knows the ground truth. The code used to produce this result is available at [9].
II-B Fixed stepsize: Strengths and weaknesses
If the unknown function $f$ is time-invariant and the stepsize $\alpha_{t}$ decreases to 0 sufficiently slowly over time (specifically, $\sum_{t}\alpha_{t}=\infty$ and $\sum_{t}\alpha_{t}^{2}=0$), then the SA iterates $y_{t}$ will converge to the root $y^{*}$ almost surely [28].
Conversely, if the step size is constant, i.e., $\alpha_{t}=\alpha$ for all $t$, and $f$ remains time-invariant, the sequence of iterates $y_{t}$ will form a Markov chain which, under suitable conditions [29], will converge exponentially fast to its stationary distribution, with a rate of convergence proportional to the step size $\alpha$. Hence, in this transient phase, higher values of $\alpha$ result in faster convergence. Once the stationary distribution is reached, the Markov chain fluctuates around the root $y^{*}$ within a region of size $\mathcal{O}(\alpha)$. Thus, lower values of $\alpha$ help reducing the iterates’ variance in the stationary phase.
When the unknown function $f$ is time-dependent, using a constant stepsize is a common design choice to effectively track the time-varying root, as in the case of OLLA. Yet, the considerations above also suggest that there is no universal stepsize value that performs well for OLLA in all scenarios:
  • When the SINR changes abruptly, a large stepsize is needed to follow the rapid root fluctuations;
  • When the SINR stabilizes, a smaller stepsize is preferable to avoid significant MCS fluctuations, which would degrade the achieved spectral efficiency. Indeed, when OLLA overshoots by selecting a high MCS, the offset is harshly penalized, causing the MCS to drop well below the optimal value. This behavior is depicted in Figure 1.
Our proposed SALAD algorithm is designed to combine the advantages of both strategies: it increases the speed of link adaptation during abrupt SINR changes, and reduces it when the SINR stabilizes, thereby ensuring accurate channel inference while mitigating inefficient MCS fluctuations.
II-C The impact of delayed HARQ feedback
The analogy between OLLA and SA holds in the absence of HARQ feedback delay, i.e., $\delta=0$. Yet, in real systems, feedback delay is unavoidable and can lead to performance degradation. For instance, suppose that the channel quality deteriorates at time $t$, but the corresponding ACK/NACKs are reported at $t+\delta_{t}$. During the interval $t,t+\delta_{t}]$, OLLA continues to select a high MCS, resulting in a series of NACKs received after $t+\delta_{t}$. This causes the SINR estimate to drop sharply, potentially well below the actual SINR value, ultimately causing the selection of over-conservative MCS.
SALAD is more robust to HARQ feedback delay, because it updates the SINR estimate using the likelihood of receiving an ACK/NACK, instead of applying a fixed additive step.
III SALAD: Self-Adaptive Link Adaptation
Our link adaptation algorithm, called SALAD (self-adaptive link adaptation), consists of three modules:
  • SINR inference (Section III-A): The SINR is estimated by binary cross-entropy loss minimization. Unlike OLLA, the SINR update is adaptive and depends on the discrepancy between the ACK/NACK observations and their expected likelihood.
  • MCS selection (Section III-B): When the SINR is deemed to be underestimated—via hypothesis testing—a high MCS value is probed. This creates a virtuous cycle, enabling faster tracking of sudden SINR surges.
  • Long-term BLER target enforcement (Section III-C): A feedback-loop mechanism enforces the BLER target $\tau$, compensating for excessive probing when necessary.
A schematic diagram of SALAD is provided in Figure 3, its pseudo-code is shown in Algorithm 3, while SALAD’s qualitative behavior is illustrated in Figure 2.
Figure 2: SALAD’s main features. High values of the bias score ratio indicate SINR underestimation, triggering the probing of high instantaneous BLER target $\tau_{t}$ and corresponding MCS indices. This results in a quick rate and SINR adaptation. The instantaneous BLER target $\tau_{t}$ is adjusted according to the integral error $E_{t}$, steering the long-term BLER towards its target $\tau$.
III-A SINR inference
We first illustrate how SALAD produces an estimate $\widetilde{\gamma}_{t}$ of the SINR at slot $t$ from past MCS values and the corresponding ACK/NACK feedback. Our method works across two different time-scales. At every slot, a simple “student” iterate infers the next SINR from the last estimate. Such iterate depends on a learning rate called $\varepsilon$. Optionally, every $N_{\varepsilon}$ slots, a powerful “teacher” model infers the whole time-series of past SINR. The student’s learning rate is finally optimized to best mimic the teacher’s behavior over the past slots. This allows SALAD to capture long-term SINR volatility patterns. We start by presenting our teacher model.
Teacher model. We introduce a parametric model $\widetilde{\gamma}_{i}(\theta)$ of the estimated SINR at each past slot $i\leq t$, parametrized by $\theta\in\mathbb{R}^{K}$. To infer the SINR evolution, we seek the parameters $\theta$ that minimize the discrepancy between the observed ACK/NACKs and the corresponding likelihood, quantified by the empirical binary cross-entropy (BCE) loss. Our “teacher” SINR inference formulation is as follows:
$\theta^{*}_{t}:=\operatorname{argmin}_{\theta}\ \sum_{i<t}\text{\small{BCE}}_{i}(\theta)+\Omega(\theta),$ (8)
where $\Omega(.)$ is a regularizer, promoting temporal smoothness of the estimated SINR, and
$\text{\small{BCE}}_{i}(\theta):=$ $\,-1{\thinspace}\mathrm{I}^{\textsc{NACK}}_{i}\log\Pr\left(1{\thinspace}\mathrm{I}^{\textsc{NACK}}_{i}=1\,|\,u_{i},\widetilde{\gamma}_{i}(\theta),b_{i}\right)$
$-\big(1$- $1{\thinspace}\mathrm{I}^{\textsc{NACK}}_{i}\big)\log\Pr\left(1{\thinspace}\mathrm{I}^{\textsc{NACK}}_{i}=0\,|\,u_{i},\widetilde{\gamma}_{i}(\theta),b_{i}\right).$ (9)
The probability of receiving a NACK in an equivalent AWGN channel can be pre-computed and stored in BLER tables $\text{\small{BLER}}^{\mathrm{tab}}$, for every MCS $u$, SINR $\gamma$ and TBS $b$. Then, (9) can be rewritten as
$\text{\small{BCE}}_{i}(\theta)=$ $\,-1{\thinspace}\mathrm{I}^{\textsc{NACK}}_{i}\log\text{\small{BLER}}^{\mathrm{tab}}_{i}(\theta)$
$-\big(1-1{\thinspace}\mathrm{I}^{\textsc{NACK}}_{i}\big)\log\left(1-\text{\small{BLER}}^{\mathrm{tab}}_{i}(\theta)\right),$ (10)
where we define:
$\text{\small{BLER}}^{\mathrm{tab}}_{i}(\theta):=\text{\small{BLER}}^{\mathrm{tab}}\left(u_{i},\widetilde{\gamma}_{i}(\theta),b_{i}\right).$ (11)
Gradient computation: To solve (8) via gradient descent, we must compute the derivative of the loss with respect to $\theta$. To obtain a convenient closed form, we approximate $1-\text{\small{BLER}}^{\mathrm{tab}}$ as a suitably translated and scaled sigmoid $\sigma$, as in [30]:
$1\!-\!\text{\small{BLER}}^{\mathrm{tab}}(u,\gamma,b)\!\approx\!\left(1+e^{-\frac{\gamma-c(u,b)}{s(u,b)}}\right)^{-1}\!\!:=\!\sigma_{u,b}(\gamma),$ (12)
where $c(u,b)$ and $s(u,b)$ are the center and scale of the sigmoid, respectively, which can be obtained via simple mean-square curve fitting. When applied on the BLER tables provided by Sionna [31], such approximation results in an average mean square loss of $4\text{\times}{10}^{-4}$, which is acceptable since our BLER target value $\tau$ is on the order of ${10}^{-1}$. Table II reports a few cases, while the full set of results is available at [9].
TABLE II: Example center and scale parameters of the sigmoid approximation (12) for the BLER for different code block size (CBS), considering PDSCH and MCS table index 2, see Section 5 in [8].
MCS index $u$ CBS (n. bits) center $c$ scale $s$
2 100 -1.91 0.44
2 2000 -2.01 0.36
6 100 4.84 0.51
6 2000 5.04 0.20
10 100 8.36 0.52
10 2000 9.04 0.04
14 100 12.20 0.57
14 2000 12.32 0.38
20 100 18.10 0.69
20 2000 18.54 0.06
By exploiting the property $\sigma^{\prime}=\frac{1}{s}\sigma(1-\sigma)$, we obtain
$\nabla_{\theta}\text{\small{BCE}}_{i}(\theta)\approx s_{i}^{-1}\left(1{\thinspace}\mathrm{I}^{\textsc{NACK}}_{i}-\text{\small{BLER}}^{\mathrm{tab}}_{i}(\theta)\right)\nabla_{\theta}\widetilde{\gamma}_{i}(\theta).$ (13)
Piece-wise linear model: A simple yet effective model is the piece-wise linear one, also called first-order spline:
$\widetilde{\gamma}_{i}(\theta)=\sum_{k=1}^{K}\widetilde{\gamma}^{k}(i,\theta_{k}),$ (14)
where $\widetilde{\gamma}^{k}(i,\theta_{k})$ is a triangle-shaped function centered at $t_{k}$ with peak height $\theta_{k}$:
$\widetilde{\gamma}^{k}(i,\theta_{k}):=\left\{\begin{array}[]{l}\theta_{k}\tfrac{t_{k+1}-i}{t_{k+1}-t_{k}}\quad\text{if }i\in[t_{k},t_{k+1}]\\ \theta_{k}\frac{i-t_{k-1}}{t_{k}-t_{k-1}}\quad\text{if }i\in[t_{k-1},t_{k}]\\ 0\quad\text{otherwise.}\end{array}\right$. (15)
In this case, a reasonable regularizer is the total variation one:
$\Omega(\theta):=\beta\sum_{k=2}^{K}\left(\theta_{k}-\theta_{k-1}\right)^{2},\quad\beta\geq 0.$ (16)
For the linear model, the gradients $\nabla_{\theta}\widetilde{\gamma}_{i}(\theta)$, $\nabla_{\theta}\Omega(\theta)$ write:
$\tfrac{\partial}{\partial\theta_{k}}\gamma_{i}(\theta)=$ $\,\left\{\begin{array}[]{l}\tfrac{t_{k+1}-i}{t_{k+1}-t_{k}}\quad\text{if }i\in[t_{k},t_{k+1}]\\ \frac{i-t_{k-1}}{t_{k}-t_{k-1}}\quad\text{if }i\in[t_{k-1},t_{k}]\\ 0\qquad\quad\ \ \text{otherwise}\end{array}\right.,\ \forall k$ (20)
$\tfrac{\partial}{\partial\theta_{k}}\Omega(\theta)=$ $\left\{\begin{array}[]{l}2\beta(2\theta_{k}-\theta_{k-1}-\theta_{k+1}),\ 1<k<K\\ 2\beta(\theta_{k}-\theta_{k+1}),\qquad\qquad k=1\\ 2\beta(\theta_{k}-\theta_{k-1}),\qquad\qquad k=K.\end{array}\right$. (24)
The values of $\alpha$ and $K$ can be determined by cross-validation.
The time-series $\{\widetilde{\gamma}_{i}(\theta^{*}_{t})\}_{i<t}$, where $\theta^{*}_{t}$ is the solution to (8), is our teacher model. The SINR estimation at slot $t$ can be extrapolated as $\widetilde{\gamma}_{t}(\theta^{*}_{t})$.
Figure 3: SALAD’s workflow diagram.
Student model. It may be computationally prohibitive to re-optimize the teacher model at every slot to produce the estimate $\widetilde{\gamma}_{t}(\theta^{*}_{t})$. We then propose a computationally cheaper but effective heuristic alternative, called student inference, which infers the SINR at slot $t$ as a gradient update of the previous estimate:
$\widetilde{\gamma}_{t}=\widetilde{\gamma}_{t-1}-\varepsilon\frac{\partial}{\partial\,\widetilde{\gamma}_{t-1}}\text{\small{BCE}}_{t-1}(\widetilde{\gamma}_{t-1})\quad\forall\,t>0$ (25)
where $\varepsilon>0$ is the learning rate and $\widetilde{\gamma}_{0}$ is initialized arbitrarily. By using the sigmoid approximation (12) and exploiting the property $\sigma^{\prime}=\frac{1}{s}\sigma(1-\sigma)$, we can rewrite (25) in closed-form:
$\widetilde{\gamma}_{t}=\widetilde{\gamma}_{t-1}+\varepsilon s_{t-1}^{-1}\left(\text{\small{BLER}}^{\mathrm{tab}}_{t-1}-1{\thinspace}\mathrm{I}^{\textsc{NACK}}_{t-1}\right)\quad\forall\,t>0,$ (26)
where we define:
$\text{\small{BLER}}^{\mathrm{tab}}_{t}:=\text{\small{BLER}}^{\mathrm{tab}}\left(u_{t},\widetilde{\gamma}_{t},b_{t}\right)\quad\forall\,t\geq 0$ (27)
and $s_{t}:=s(u_{t},b_{t})$ is the sigmoid scale corresponding to MCS $u_{t}$ and TBS $b_{t}$.
Surprise effect: We observe that the student’s SINR increment in (26) is proportional to the term $\text{\small{BLER}}^{\mathrm{tab}}_{t-1}-1{\thinspace}\mathrm{I}^{\textsc{NACK}}_{t-1}$ which quantifies the “surprise” of observing $1{\thinspace}\mathrm{I}^{\textsc{NACK}}_{t-1}$ relative to its expected likelihood $\text{\small{BLER}}^{\mathrm{tab}}_{t-1}$. Indeed,
  • Suppose $\text{\small{BLER}}^{\mathrm{tab}}_{t-1}\approx 1$:
    • If an ACK is unexpectedly received ($1{\thinspace}\mathrm{I}^{\textsc{NACK}}_{t-1}=0$), then $\widetilde{\gamma}$ is increased significantly, confirming that the SINR is currently underestimated.
    • Receiving a NACK aligns with expectations and has virtually no impact on the SINR update;
  • Conversely, if $\text{\small{BLER}}^{\mathrm{tab}}_{t-1}\approx 0$:
    • A NACK indicates SINR overestimation, resulting in a significant SINR decrease;
    • an ACK only slightly affects the SINR estimation.
Knowledge distillation. For the student SINR estimate (26) to accurately track actual SINR variations, the learning rate $\varepsilon$ should ideally increase with the volatility of the SINR. Yet, the ground truth is inherently unobservable. To overcome this, we adopt a procedure inspired by knowledge distillation in neural networks [32], where the student estimator learns to imitate a more powerful teacher estimator. Specifically, the teacher SINR estimator is first optimized by solving (8) according to recent observations every $N_{\varepsilon}$ slots. Then, the student estimator (26) is replayed on the same data using different values of $\varepsilon$, and the value that yields the closest match to the reference estimator is selected for use in the subsequent slots. This procedure is described in Algorithm 2.
Algorithm 2 Knowledge Distillation to Adapt Learning Rate
1:At slot $t$, optimize teacher model (8). Obtain $\widetilde{\gamma}_{i}(\theta^{*}),\,\forall i\!\leq\!t$
2:Replay (26), for different $\varepsilon$; obtain $\widetilde{\gamma}^{(\varepsilon)}_{i},\ \forall\,i\leq t$ and $\forall\,\epsilon$
3:Select student’s learning rate as
$\varepsilon^{*}=\operatorname{argmin}_{\varepsilon}\sum_{i<t}\left(\widetilde{\gamma}_{i}(\theta^{*})-\widetilde{\gamma}_{i}^{(\varepsilon)}\right)^{2}.$
Remark: SALAD’s SINR inference is agnostic to CQI reports. Yet, one can incorporate them as in OLLA, by interpreting $\widetilde{\gamma}$ as an offset to be applied to the most recent CQI report.
III-B MCS selection
SALAD selects the MCS $u_{t}$ at slot $t$ via ILLA under an optimized instantaneous BLER target $\tau_{t}$, i.e.,
$u_{t}:=u^{\mathrm{ILLA}}(\widetilde{\gamma}_{t-\delta_{t}},\tau_{t},b_{t})\quad\forall\,t\geq 0.$ (28)
Our dynamic target selection philosophy is the following:
  • If the SINR inference is considered accurate at time $t$, we select the instantaneous target optimally as $\tau_{t}=\tau$.
  • If the SINR is detected to be underestimated, a high instantaneous BLER target $\tau_{t}:=\tau^{\mathrm{probe}}\gg\tau$ is probed. This corresponds to a high MCS index and a high BLER ($\text{\small{BLER}}^{\mathrm{tab}}_{t}\approx 1$). Receiving an ACK ($1{\thinspace}\mathrm{I}^{\textsc{NACK}}_{t}=0$) in this setting provides two advantages:
    • i)
      The spectral efficiency immediately increases;
    • ii)
      Thanks to the “surprise effect”, the inferred SINR quickly catches up with the ground truth. This creates a virtuous circle that improves the accuracy of future MCS selections.
We now turn to the design of a method for detecting SINR underestimation.
Detection of SINR underestimation and MCS probing. The quality of SINR estimation can be assessed by comparing the BLER of recent transmissions (estimated from pre-computed tables under the assumption of accurate SINR) and the corresponding ACK/NACKs, providing an empirical BLER estimate. Intuitively, if the BLER significantly exceeds the observed portion of NACKs, the former is being overestimated, while the SINR is underestimated. To formalize this, we use a simple statistical hypothesis test.
We first introduce the calibration bias score, summing up all signed differences between the BLER values and the ACK/NACK observations $1{\thinspace}\mathrm{I}^{\textsc{NACK}}$ over the last $T$ samples:
$\mathcal{S}_{t}=\frac{1}{T}\sum_{i=1}^{T}\text{\small{BLER}}^{\mathrm{tab}}_{t-\delta_{t}-i}-1{\thinspace}\mathrm{I}^{\textsc{NACK}}_{t-\delta_{t}-i}.$ (29)
The score has zero mean under the null hypothesis $\mathcal{H}_{0,t}$ that all BLER and SINR estimates are correct, where
$\mathcal{H}_{0,t}:=\!\left\{\mathbb{E}[1{\thinspace}\mathrm{I}^{\textsc{NACK}}_{i-\delta_{t}}]\!=\!\text{\small{BLER}}^{\mathrm{tab}}_{i-\delta_{t}},\ t-T\leq i\leq t-1\right\}.$ (30)
We reject the null hypothesis $\mathcal{H}_{0}$, and conclude that the SINR estimation is poor, when the value of $\mathcal{S}_{t}$ is unlikely under $\mathcal{H}_{0}$. More specifically, we deem the SINR underestimated when $\mathcal{S}_{t}$, normalized by its variance under $\mathcal{H}_{0}$, exceeds a design parameter $\rho>0$:
$\frac{\mathcal{S}_{t}}{\sqrt{\mathrm{Var}[\mathcal{S}_{t}|\mathcal{H}_{0,t}]}}>\rho\ \$$\ \ \Rightarrow\ \ \text{SINR is underestimated,}$ (31)
where
$\mathrm{Var}\left[\mathcal{S}_{t}|\mathcal{H}_{0,t}\right]=\frac{1}{T^{2}}\sum_{i=1}^{T}\text{\small{BLER}}^{\mathrm{tab}}_{t-\delta_{t}-i}(1-\text{\small{BLER}}^{\mathrm{tab}}_{t-\delta_{t}-i}).$ (32)
In this case, as motivated above, a high BLER target $\tau_{t}:=\tau^{\mathrm{probe}}\gg\tau$ is selected, resulting in the probing of an aggressive MCS value.
Remark. A non-negligible HARQ feedback delay may render the score $\mathcal{S}_{t}$ temporarily inaccurate and trigger a sequence of poor probing decisions and a transient BLER increase. To mitigate this behavior, we recommend probing high MCS values only with some probability $p^{\mathrm{probe}}$ if (31) holds.
III-C Long-term BLER target enforcement
An overly aggressive MCS probing strategy as defined in (31) can lead to a transient increase in BLER. Moreover, according to (26), the inferred SINR is only mildly decreased following a NACK, which may promote further probing in future transmissions. To address this, we introduce a feedback mechanism that adapts the instantaneous BLER target $\tau_{t}$ based on the achieved BLER, inspired by the Proportional-Integral-Derivative (PID) control approach in [33].
Specifically, we define the integral error
$E_{t}:=\sum_{i<t-\delta_{t}}\tau-1{\thinspace}\mathrm{I}^{\textsc{NACK}}_{i}$ (33)
and adapt the instantaneous BLER target $\tau_{t}$ as follows:
$\tau_{t}\leftarrow\mathrm{clip}\left(\tau_{t}+k_{E}E_{t},0,1\right),\quad k_{E}>0.$ (34)
Observe that if the portion of NACKs exceeds the BLER target $\tau$, $E_{t}$ becomes negative, resulting in a lower BLER target and a more conservative MCS selection. Conversely, a positive $E_{t}$ leads to a more aggressive MCS selection. This mechanism is instrumental for maintaining the BLER, computed over the whole connection duration, close to the desired target $\tau$.
The full SALAD’s pseudo-code is outlined in Algorithm 3.
Algorithm 3 SALAD: Self-Adaptive Link Adaptation
1:Parameters $k_{E}>0$, $\rho>0$, $p^{\mathrm{probe}}\in(0,1)$, $\tau^{\mathrm{probe}}>\tau$.
2:Initialize $E_{I}=0$, $\widetilde{\gamma}$ arbitrarily.
3:for slot $t=0,1,\dots$ do
4:  Receive set $\mathcal{N}$ of ACK/NACK. Let $\mathcal{U},\mathcal{B}$ be the corresponding MCS and TBS, respectively.
5:  for $1{\thinspace}\mathrm{I}^{\textsc{NACK}}\in\mathcal{N}$ do
6:   Update integral error: $E\leftarrow E+\tau-1{\thinspace}\mathrm{I}^{\textsc{NACK}}$
7:  end for
8:# Student’s SINR inference update
9:  for $(1{\thinspace}\mathrm{I}^{\textsc{NACK}},u,b)\in(\mathcal{N},\mathcal{U},\mathcal{B})$ do
$\widetilde{\gamma}\leftarrow\widetilde{\gamma}+\varepsilon s^{-1}\left(\text{\small{BLER}}^{\mathrm{tab}}(u,\widetilde{\gamma},b)-1{\thinspace}\mathrm{I}^{\textsc{NACK}}\right)$
10:  end for
11:  Set the latest SINR estimate to $\widetilde{\gamma}_{t-\delta_{t}}:=\widetilde{\gamma}$ [ $\text{\,}\mathrm{dB}\text{/}$]
12:  if user is scheduled then
13:    # SINR underestimation detection
14:   Compute bias score $\mathcal{S}_{t}$ and its variance $\mathrm{Var}[\mathcal{S}_{t}|\mathcal{H}_{0,t}]$ as in (29), (32), respectively.
15:   if $\tfrac{\mathcal{S}_{t}}{\sqrt{\mathrm{Var}[\mathcal{S}_{t}|\mathcal{H}_{0}]}}>\rho\land\mathrm{rand}(0,1)<p^{\mathrm{probe}}$ then
16:    # MCS probing
17:     Set $\tau_{t}:=\tau^{\mathrm{probe}}$
18:   else
19:     Set $\tau_{t}:=\tau$
20:   end if
21:   # BLER target $\tau$ enforcement
22:   Adjust $\tau_{t}\leftarrow\mathrm{clip}\left(\tau_{t}\!+\!k_{E}E,0,1\right)$
23:    # MCS selection
24:   Assign MCS $u_{t}=u^{\mathrm{ILLA}}(\widetilde{\gamma}_{t-\delta_{t}},\tau_{t},b_{t})$
25:  end if
26:  if $\mod(t,N_{\varepsilon})=0$ then
27:    # Knowledge distillation
28:   Update the learning rate $\varepsilon$ via Algorithm 2
29:  end if
30:end for
31:Return $u_{t}$
III-D A practical consideration
For large CBS values, the BLER exhibits a steep transition between its extreme values, 1 and 0, near the Shannon SINR limit. This is illustrated in Table II, showing smaller scale parameter $s$ for larger CBS, indicating a steeper BLER cutoff.
As a result, the term $s_{t-1}^{-1}\left(\text{\small{BLER}}^{\mathrm{tab}}_{t-1}-1{\thinspace}\mathrm{I}^{\textsc{NACK}}_{t-1}\right)$, which appears in SALAD’s SINR update (26), can take on extreme values when the TBS is large. Specifically, upon probing, a received ACK leads to a substantial SINR increase $\widetilde{\gamma}_{t}\gg\widetilde{\gamma}_{t-1}$. Conversely, receiving a NACK results in only a negligible update $\widetilde{\gamma}_{t+1}\approx\widetilde{\gamma}_{t}$. In the absence of probing, the situation is reversed: an ACK leads to a negligible SINR update, while a NACK causes a significant SINR decrease.
A simple and effective remedy is to clip the values of $\text{\small{BLER}}^{\mathrm{tab}}_{t}$ and $s_{t}$ to appropriate intervals, e.g., $(0.01,0.99)$ and $(0.5,10)$, respectively.
III-E The impact of delayed HARQ feedback
We conclude the presentation of SALAD by analyzing the impact of delayed HARQ on SINR inference and MCS selection, while emphasizing its connection to OLLA.
Absence of HARQ feedback delay ($\delta=0$). In this case, SALAD’s SINR inference update can be interpreted as a variant of OLLA with adaptive instantaneous BLER target and offset parameter $\Delta^{\textsc{NACK}}$. In fact, since the MCS is selected via the up-to-date SINR estimation, i.e., $u_{t-1}=u^{\mathrm{ILLA}}_{t-1}(\widetilde{\gamma}_{t-1},\tau_{t})$, we can approximate the target $\tau_{t-1}$ as the effective BLER:
$\tau_{t-1}\approx\text{\small{BLER}}^{\mathrm{tab}}_{t-1}:=\text{\small{BLER}}^{\mathrm{tab}}(u_{t-1},\widetilde{\gamma}_{t-1}),\ \text{if }\delta=0.$ (35)
Then, SALAD’s SINR update (26) can be rewritten as
$\widetilde{\gamma}_{t}\approx\widetilde{\gamma}_{t-1}+\varepsilon s_{t-1}^{-1}\left(\tau_{t-1}-1{\thinspace}\mathrm{I}^{\textsc{NACK}}_{t-1}\right)\quad\forall\,t>0$ (36)
which corresponds to OLLA’s update—cfr. with (7):
$\widetilde{\gamma}^{\mathrm{OLLA}}_{t}=\widetilde{\gamma}^{\mathrm{OLLA}}_{t-1}+\frac{\Delta^{\textsc{NACK}}_{t-1}}{1-\tau_{t-1}}\left(\tau_{t-1}-1{\thinspace}\mathrm{I}^{\textsc{NACK}}_{t-1}\right)$ (37)
with a time-adaptive offset parameter $\Delta^{\textsc{NACK}}_{t-1}:=\varepsilon s^{-1}_{t-1}(1-\tau_{t-1})$ and a dynamic target $\tau_{t-1}$.
Presence of HARQ feedback delay ($\delta>0$). In this scenario, SALAD’s SINR update differs significantly from that of time-adaptive OLLA. In both algorithms, the MCS $u_{t-1}$ is selected based on an outdated SINR estimation, specifically, $u_{t-1}=u^{\mathrm{ILLA}}_{t-1}(\widetilde{\gamma}_{t-1-\delta_{t-1}},\tau_{t-1})$. Yet, the SINR staleness renders (35) invalid: the target $\tau_{t-1}$, used by OLLA in its SINR update, is different from the effective BLER, $\text{\small{BLER}}^{\mathrm{tab}}_{t-1}$, used by SALAD. In other words, SALAD updates the SINR using fresher information than OLLA. As a result, SALAD’s SINR estimation is more robust to feedback delay, as confirmed by the experiments in Section IV-A.
IV Experimental Results
To evaluate SALAD’s performance, we conduct full-stack system simulations and real-world over-the-air (OTA) measurements using a real-time 5G testbed. Specifically, we implement SALAD [^5] A slight variation of SALAD is tested, where the instantaneous BLER target is adjusted via integral error (line 21) only if probing is not performed. and OLLA in the 5G downlink (PDSCH) scheduler within the gNB multiple access (MAC) layer of the OpenAirInterface (OAI) [34] using the Sionna Research Kit (SRK) [35]. Using OAI’s radio frequency simulator (RFSim), we perform full-stack simulations aligned with the 5G NR standard. Furthermore, leveraging the NVIDIA Aerial CUDA-accelerated RAN platform, we carry out real-world OTA measurements with a commercial off-the-shelf (COTS) Open RAN radio unit (O-RU) and a COTS user equipment (UE) deployed inside an anechoic chamber. All experiments in this section adopt a long-term BLER target value of $\tau=0.1$, a widely used setting in the literature [2, 3].
TABLE III: SALAD’s configuration parameters. Default parameters are designed manually. The other two configurations are optimized for a specific low/high throughput regime by Nelder-Mead method.
Parameter Default (manual design) Optimized for low throughput Optimized for high throughput
$\varepsilon$ 1 1.09 0.85
$\rho$ 0.25 0.335 0.346
$T$ 15 13 13
$p^{\mathrm{probe}}$ 0.15 0.134 0.139
$\tau^{\mathrm{probe}}$ 0.999 0.941 0.957
$k_{E}$ 0.01 0.032 0.0017
(a) Channel gain surge: No CQI feedback, SISO with 1 O-RU antenna, 24 PRBs.
(b) Channel gain drop: With CQI feedback, single-layer MISO with 2 O-RU antennas, 51 PRBs.
Figure 4: Controlled SINR. SINR estimation and MCS index for SALAD and OLLA from OAI RFSim with 7DS2U TDD slot pattern.
IV-A 5G testbed with controlled SINR
We first evaluate SALAD’s performance via full-stack, system-level simulations using OAI’s RF simulator (RFSim). The pathloss in RFSim is controlled via a testbench script that alters the channel gain at predefined timestamps. The script also loads application-layer traffic from an external traffic generator server (iperf3 -s) to a traffic generator app (iperf3) running on OAI’s software-defined UE, which is connected to the software-defined gNB via RFSim. We simulate two RF channel configurations: (i) a single-input single-output (SISO) AWGN channel without CQI feedback, i.e., CSI-RS was disabled, and (ii) a multiple-input single-output (MISO) AWGN channel with CQI feedback. For the frame structure, we use the 5G NR N78 band with a 7DS2U (7 DL slots, 1 special slot, 2 UL slots) TDD slot pattern. RFSim does not run in real-time and advances according to simulation speed. Yet, higher protocol layers, e.g., MAC and above, run in real system time and are 5G-NR-compliant. SALAD’s default parameters (see Table III) are used.
Channel surge. Figure 4(a) shows the time evolution of the scheduled MCS indices for SALAD and OLLA (top), along with their corresponding SINR estimates and the channel gain (bottom). The channel gain surges from $3\text{\,}\mathrm{dB}\text{/}$ to $7\text{\,}\mathrm{dB}\text{/}$ at timestamp $t=0$. The UDP downlink throughput is $1\text{\,}\mathrm{Mbit}\text{/}\mathrm{s}$. Figure 4(a) shows that SALAD adapts to the improved channel conditions approximately twice as fast as OLLA (with $\Delta^{\text{NACK}}=1$), thanks to better SINR inference enabled by MCS probing. In fact, upon the channel gain surge, SALAD probes high MCS indices (top) which translate to better SINR inference (bottom). However, SALAD also probes high MCS values when unnecessary (top), particularly before the surge. This can be mitigated by decreasing the probing probability $p^{\mathrm{probe}}$ and/or increasing the calibration ratio $\rho$.
Channel drop, delayed feedback. We enable CSI-RS and use a variant of SALAD that incorporates CQI reports $\gamma^{\mathrm{rep}}$ by simply considering $\gamma^{\mathrm{rep}}+\widetilde{\gamma}$ as the current SINR estimate, similarly to OLLA. Figure 4(b) shows the time evolution of the scheduled MCS indices of SALAD and OLLA (solid lines, top) along with the corresponding MCS derived from the reported CQI ($\mathrm{MCS}_{\mathrm{CQI}}$) [8], as well as their SINR estimates and the channel gain (bottom). The channel gain drops from $13\text{\,}\mathrm{dB}\text{/}$ to $3\text{\,}\mathrm{dB}\text{/}$ at time $t=0$. As before, the DL UDP throughput is $1\text{\,}\mathrm{Mbit}\text{/}\mathrm{s}$. We observe from Figure 4(b) that, immediately after the channel degradation, SALAD and OLLA (with $\Delta^{\text{NACK}}=1$) decrease the MCS at a similar rate. However, the delayed CQI report, received at time $t\approx\,$$0.3\text{\,}\sec$, causes the estimated SINR to drop further, resulting in an undershoot in both SINR and MCS for both algorithms. SALAD promptly compensates for this undershoot by resuming the probing of higher MCS values, enabling a faster recovery compared to OLLA.
This behavior stems from Section III-E: SALAD is robust to feedback inaccuracies since it updates its SINR according to the effective BLER which, under delayed feedback, may considerably differ from the BLER target considered by OLLA.
IV-B Over-the-air experiments
Next, we benchmark the performance of SALAD against that of OLLA in a real-time over-the-air (OTA) 5G testbed. Our implementation leverages the Aerial RAN CoLab Over-the-Air (ARC-OTA) platform [36], which integrates NVIDIA’s Aerial CUDA-Accelerated RAN (using NVIDIA cuPHY Layer 1 [37]) with OAI’s higher protocol stacks [34].
Figure 5 shows our controllable radio environment inside an anechoic chamber. Two UEs are placed in the left corner, one behind and one above a rotational table. A flat RF absorber panel is mounted on the opposite side of the rotational table, facing the UEs. The table is shown in Figure 5 at a reference rotation angle of $\phi=$280\text{\,}\mathrm{\SIUnitSymbolDegree}\text{/}$\$. Depending on the table’s rotation angle, the UEs experience either line-of-sight (LoS) or non-line-of-sight (NLoS) conditions with respect to the O-RU. In the right corner, we place a COTS O-RU with 4 antennas and operating with a $100\text{\,}\mathrm{MHz}\text{/}$ channel bandwidth (273 PRBs) in the 5G N78 band. The O-RU’s transmit power is set to $9\text{\,}\mathrm{dBm}\text{/}$ per antenna to prevent MCS saturation in LoS while maintaining connectivity in NLoS.
Experimental setup. A script controls the table rotation while recording the rotation angles and the corresponding timestamps. The table is equipped with a high-precision rotation axis motor which is re-calibrated after each power cycle. This setup ensures that the RF channel characteristics remain consistent throughout our measurement campaign. We use a COTS UE (OnePlus Nord) and rotate the rotational table back and forth. The rotation begins in a LoS position ($\phi=$100\text{\,}\mathrm{\SIUnitSymbolDegree}\text{/}$\$, i.e., the absorber’s location is opposite to its placement in Figure 5) and proceeds at constant angular speed until the UE is in NLoS position ($\phi=$190\text{\,}\mathrm{\SIUnitSymbolDegree}\text{/}$\$, with the absorber closest to the UE). Then, the table rotates back to the starting position. The rotation speed is set to $1300\text{\,}\mathrm{\SIUnitSymbolDegree}\text{/}\mathrm{min}$, resulting in a full back-and-forth cycle lasting $8\text{\,}\sec$. CQI reports are observed to be inaccurate and unreliable. For this reason, CQI reports are disabled for our OTA experiments.
Figure 5: Measurement setup in anechoic chamber at ETH Zürich with COTS UE and rotational absorber table (left) and COTS O-RU (right).
Default SALAD vs. default OLLA. Figure 6 provides a first visual comparison of the performance of SALAD and OLLA in OTA experiments, shown as a function of rotation angle. SALAD uses its manually designed “default” configuration (see Table III), while OLLA employs an offset update $\Delta^{\text{NACK}}=1$, which is regarded as the default value in the literature, see [17, 22]. All experiments are obtained under identical controlled RF conditions with a TCP throughput reaching $180\text{\,}\mathrm{Mbit}\text{/}\mathrm{s}$ in LoS and $70\text{\,}\mathrm{Mbit}\text{/}\mathrm{s}$ in NLoS.
Figure 6 superposes the results from 30 rotation trajectories. The individual trajectories are then overlayed and shown in light shading. The median of all trajectory data falling within the same angle bin is plotted in dark thick lines. For clarity of visualization, Figure 6 focuses exclusively on the backward trajectory, i.e., from NLoS to LoS.
The upper subplot in Figure 6 shows the MCS selected by SALAD and OLLA. SALAD consistently selects higher MCS values and better tracks the channel gain surge, whereas OLLA tends to oscillate more under stable channel conditions, resulting in efficiency loss with respect to SALAD. The second subplot in Figure 6 shows the “normalized” throughput (TP)
$\text{Normalized TP}=\frac{\text{TP}_{\text{1st-round}}\times\text{SE}_{\text{sched}}}{\text{SE}_{\text{min}}},$ (38)
defined as the throughput achieved during the first HARQ round ($\text{TP}_{\text{1st-round}}$), i.e., computed only for packets which do not require retransmission, multiplied by the SE of the scheduled MCS ($\text{SE}_{\text{sched}}$) and scaled by the minimum available SE ($\text{SE}_{\text{min}}$). This performance metric jointly captures SE, raw connection speed, and reliability. It is more practically relevant than the classic expected SE formulation in (1). SALAD consistently achieves higher normalized throughput than OLLA with gains (in orange) reaching up to 30%. Indeed, selecting a higher MCS and achieving a higher instantaneous rate can enable the TCP congestion control mechanism to further increase its future transmission rate, thereby creating a virtuous circle.
Although not visualized here, both SALAD and OLLA can maintain the long-term BLER, computed over the entire connection duration, at the target value $\tau=0.1$. The third subplot in Figure 6 presents the short-term BLER, computed over a sliding window of approximately $250\text{\,}\mathrm{ms}\text{/}$. For both algorithms, the short-term BLER fluctuates around the target $\tau=0.1$. SALAD exhibits a slight BLER over-shoot (approximately $0.135$ versus $\tau=0.1$) during the channel gain surge, at an angle range of $170\text{\,}\mathrm{\SIUnitSymbolDegree}\text{/}-$$175\text{\,}\mathrm{\SIUnitSymbolDegree}\text{/}$, due to aggressive MCS probing. In contrast, OLLA’s BLER undershoots during the same interval, indicating slower adaptation to improving channel conditions. The second subplot shows that SALAD’s BLER overshoot is compensated by improved overall performance.
The lower subplot in Figure 6 shows the SINR $\widetilde{\gamma}$ estimated by SALAD and OLLA. SALAD’s SINR estimate is consistently higher and more stable than OLLA’s. Interestingly, even when the short-term BLER overshoots, the SINR estimation remains unaffected by the received NACKs. This stems directly from expression (26): when the predicted BLER is high and a NACK is received, the estimated SINR is only slightly decreased. In contrast, OLLA always penalizes NACKs with a fixed SINR decrement of $\Delta^{\textsc{NACK}}$.
Figure 6: OTA—Default SALAD and OLLA: Default SALAD (see Table III) vs. default OLLA ($\Delta^{\textsc{NACK}}=1$) in improving channel conditions (NLoS to LoS) with maximum TCP throughput, single-layer MIMO, 4 O-RU antennas, 273 PRBs, 3DSU TDD slot pattern.
(a.1) $1\text{\,}\mathrm{Mbit}\text{/}\mathrm{s}$
[…]
(a.2) $10\text{\,}\mathrm{Mbit}\text{/}\mathrm{s}$
(a) (a.3) $100\text{\,}\mathrm{Mbit}\text{/}\mathrm{s}$
(a) Normalized throughput gain vs. OLLA ($\Delta^{\textsc{NACK}}=1$).
(b.1) $1\text{\,}\mathrm{Mbit}\text{/}\mathrm{s}$
(b.2) $10\text{\,}\mathrm{Mbit}\text{/}\mathrm{s}$
(b.3) $100\text{\,}\mathrm{Mbit}\text{/}\mathrm{s}$
(b) Short-term BLER deviation from long-term target $\tau$.
Figure 8: OTA—Variants of SALAD and OLLA: Performance of various OLLA variants and SALAD with default and optimized parameters, comparing normalized TP (a) and standard error to the BLER target (b) for three different application-layer TP scenarios: $1\text{\,}\mathrm{Mbit}\text{/}\mathrm{s}$ UDP, $10\text{\,}\mathrm{Mbit}\text{/}\mathrm{s}$ UDP, and $100\text{\,}\mathrm{Mbit}\text{/}\mathrm{s}$ TCP.
One SALAD vs. many OLLAs. One may wonder if OLLA can match SALAD’s performance with proper tuning. Figure 8 shows that a single SALAD configuration can outperform or match each OLLA instance across a range of scenarios. In other words, OLLA would require proper dynamic parameter tuning to approach SALAD’s performance, while SALAD can perform well out-of-the-box in different scenarios. Specifically, we consider three traffic regimes: (i) constant $1\text{\,}\mathrm{Mbit}\text{/}\mathrm{s}$ UDP traffic, (ii) constant $10\text{\,}\mathrm{Mbit}\text{/}\mathrm{s}$ UDP traffic, and (iii) highest possible TCP throughput, averaging $100\text{\,}\mathrm{Mbit}\text{/}\mathrm{s}$. For each traffic regime, we report the achieved normalized throughput (top subplot) relative to default OLLA ($\Delta^{\textsc{NACK}}=1$) and the root mean square deviation of the instantaneous BLER from the target $\tau=0.1$ (bottom), expressed as a percentage of the target $\tau$. A lower value of the metric indicates that the BLER stays closer to the target at all times, i.e., the algorithm can promptly adapt to varying channel conditions. Both metrics are computed as trajectory mean over the full rotation cycle.
To interpret Figure 8, we first examine the manually designed “default” SALAD configuration (see Table III). Default SALAD is compared with three OLLA variants: “slow” OLLA with NACK offset $\Delta^{\textsc{NACK}}=0.5$, default OLLA ($\Delta^{\textsc{NACK}}=1$), and “fast” OLLA ($\Delta^{\textsc{NACK}}=2$). In the low-throughput regime ($1\text{\,}\mathrm{Mbit}\text{/}\mathrm{s}$), all algorithms achieve similar normalized throughput, but “slow” OLLA fails to meet the BLER target. In fact, when the throughput is low, the frequency of HARQ reports is also low, making channel variations appear rapid from the perspective of a link adaptation algorithm that updates at every HARQ feedback. This renders “slow” OLLA unable to keep up with channel variations. In the mid-throughput regime ($10\text{\,}\mathrm{Mbit}\text{/}\mathrm{s}$), all algorithms can closely meet the BLER target. Yet, in terms of throughput, SALAD performs on par with “slow” OLLA and outperforms both default and “fast” OLLA by 8.8% and an additional 15.2%, respectively. In the high-throughput regime ($100\text{\,}\mathrm{Mbit}\text{/}\mathrm{s}$), default SALAD again matches “slow” OLLA and outperforms the other OLLA variants, achieving throughput gains of 14.4% over default OLLA and an additional 25.5% over “fast” OLLA. In fact, the high frequency of HARQ feedback makes channel variations appear slow, favoring more conservative adaptation.
As a result, SALAD rightly earns the “self-adaptation” label, since it exhibits customized behavior tailored to each situation, without any dynamic parameter tuning.
Figure 9: Knowledge distillation for learning rate $\varepsilon$ self-adaptation, with $N_{\varepsilon}=200$, $\beta=0$. The number of knots $K$ is selected via cross-validation. The SINR [ $\text{\,}\mathrm{dB}\text{/}$] evolves according to a sinusoid with time-varying frequency. The learning rate $\varepsilon$ is optimized over a grid of step 0.1. As the SINR frequency increases, $\varepsilon$ also increases, allowing SALAD to track the ground truth. Results are reproducible via [9].
SALAD’s parameter tuning. Finally, we investigate the extent to which SALAD’s parameters can be customized in a specific scenario, and how this may negatively impact SALAD’s performance in different traffic regimes. To this end, SALAD’s parameters are optimized in the low throughput scenario (denoted as “opt. low TP” in Figure 8) and the high throughput scenario (“opt. high TP”) separately.
For parameter optimization we use the Nelder-Mead algorithm [38], which (i) does not require any gradient information and (ii) is well-suited for objective functions that are expensive to evaluate [39]. Such features suit our needs: (i) the objective, defined as a linear combination of normalized throughput and the negative mean squared deviation of the BLER from the target $\tau$, is not available in closed-form. Furthermore, (ii) evaluating the objective for one parameter configuration requires multiple full table rotations, lasting several tens of seconds overall, with an additional settling time of a few seconds between consecutive trials to allow transient effects to vanish. In each traffic regime, the initial configuration is set to the “default” one, and Nelder-Mead iterates over 25 different steps. The same process is repeated three times in the low throughput regime and five times in high-throughput. The resulting optimal parameters are reported in Table III.
Finally, we evaluate the performance of both fine-tuned configuration across all traffic regimes (low, mid and high throughput), as shown in Figure 8. As expected, the fine-tuned SALAD configuration outperforms the other SALAD’s variants in its preferred regime, although the gains remains limited. Moreover, the default parameters, manually selected prior to the optimization campaign, turns out to be a reasonable choice. This demonstrates that SALAD’s performance is quite stable across a range of parameter settings.
IV-C Knowledge distillation for learning rate self-adaptation
To conclude, we evaluate learning rate $\varepsilon$ online optimization via knowledge distillation (see Section III-A) under controlled SINR evolution in pure simulation. Every $N_{\varepsilon}$ slots, the history of ACK/NACK feedback, TBS, and selected MCS is collected. The “teacher” inference model—the piece-wise linear model (14)—is optimized using batch gradient descent, using the gradient expressions (13),(20). The number of knots $K$ is selected via cross-validation, by splitting the history into training and test sets, and minimizing the loss on the latter. Next, the “student” gradient-based inference algorithm is replayed on the same data using different candidate values of $\varepsilon$. The learning rate that minimizes the squared difference between the student’s and the teacher’s inferences is selected for use in the next slots. Figure 9 illustrates the qualitative effect of knowledge distillation when the SINR oscillates according to a sinusoid with time-varying frequency. The learning rate $\varepsilon$ adapts to the frequency of the unknown SINR, allowing SALAD to closely track the ground truth. As the frequency increases, $\varepsilon$ increases accordingly to follow faster SINR variations. Conversely, as the frequency decreases, $\varepsilon$ is reduced to prevent excessive fluctuations.
The code used to produce Figures 1, 2, 9 is available at [9].
V Conclusions
In this paper, we have introduced SALAD, a novel algorithm for adaptive MCS selection for wireless communications. SALAD infers the user’s effective SINR from ACK/NACK feedback alone by minimizing a regularized cross-entropy loss with respect to predicted BLER values using gradient descent. Its learning rate can be optimized at run-time via knowledge distillation, by occasionally comparing its performance against a more powerful “teacher” inference performed in hindsight. To accelerate MCS adaptation under improving channel conditions, SALAD leverages hypothesis testing to detect SINR underestimation and safely probes high MCS values. This improves both short-term spectral efficiency and long-term SINR inference accuracy. A feedback control loop ensures that the long-term BLER remains close to the desired target.
We have evaluated SALAD in OTA experiments and benchmarked it against the industry-standard OLLA. Results demonstrated that SALAD delivers robust and consistent performance across a wide range of traffic regimes, achieving up to 15% gains in throughput and spectral efficiency over default OLLA configurations while reliably meeting the BLER target. Additionally, SALAD can be fine-tuned to specific traffic regimes using derivative-free optimization techniques.
A natural extension of this work involves learning an optimal MCS selection rule conditioned on the latest SINR estimate, bias score, and integral error, thereby replacing the heuristic in lines 1524 of Algorithm 3. This approach may be extended to handle CQI reports more effectively.